# Beyond Slater’s Condition in Online CMDPs with Stochastic and Adversarial Constraints

### v1

@@ 2.3 Baseline for the Adversarial Setting
where the unconstrained optimal value is defined as $\textsc{OPT}\coloneqq\max_{q\in\Delta(M)}\frac{1}{T}\sum_{t=1}^{T}r_{t}^{\top}q$. In order to quantify $\alpha$, we introduce a problem-specific parameter $\rho\in\left[0,1\right]$, which is defined as
@@ 3 Algorithm
2:Initialize occupancy $\widehat{q}_{1}\leftarrow\frac{1}{|X_{k}||A||X_{k+1}|}$, the estimated transitions space $\mathcal{P}_{1}$ as the set of all the possible transition functions, and counters $N_{1}(x,a)=M_{1}(x^{\prime}|x,a)=0$ for all $k\in[0,...,L-1]$ and $(x,a,x^{\prime})\in X_{k}\times A\times X_{k+1}$
@@ 3.1 Initialization and Loss Estimation
Algorithm 1 receives as input the time horizon $T$, the set of states $X$, the set of actions $A$, the learning rate $\eta$, the implicit exploration factor $\gamma$, and the confidence $\delta\in(0,1)$. The occupancy measure $\widehat{q}_{1}$ is initialized uniformly over all tuples $(x_{k},a,x_{k+1})\in X_{k}\times A\times X_{k+1}$ for each layer $k\in[0,\dots,L-1]$. The transition function confidence set $\mathcal{P}_{1}$ is initialized as the set of all the possible transition functions. The counters $N_{t}(x,a)$ and $M_{t}(x^{\prime}|x,a)$, which are respectively defined as $N_{t}(x,a)=\sum_{\tau=1}^{t-1}\mathbb{I}_{\tau}\{x,a\}$ for all $(x,a)\in X\times A$, $M_{t}(x^{\prime}|x,a)=\sum_{\tau=1}^{t-1}\mathbb{I}_{t}\{x,a,x^{\prime}\}$ for all $(x,a,x^{\prime})\in X_{k}\times A\times X_{k+1},k\in[0,...,L-1],$ are initialized to 0 (see Line 2). We denote by $\mathbb{I}_{t}\{x,a\}$ and $\mathbb{I}_{t}\{x,a,x^{\prime}\}$ the indicator functions for the state-action(-state) visit at episode $\tau$.
@@ 3.2 Weights Estimation
At each episode, the counters are updated given the trajectory observed as feedback, namely, $N_{t}(x,a)$ and $M_{t}(x^{\prime}|x,a)$ are updated by incrementing by $1$ the entries of the tuples visited during the current episode. Then, the algorithm sets the weights that will be used to build the constraint estimates (Line 8). Specifically, given a pair $(x,a)\in X\times A$, $i\in[m]$, and $t\in[T-1]$, the weights $w_{t,x,a,i}$ are defined as follows:
$w_{t,x,a,i}(\tau)\coloneqq\beta_{\tau,i}(x,a)\prod_{h\in\mathcal{T}_{t-1,x,a}:h>\tau}(1-\beta_{h,i}(x,a))\quad\forall\tau\in\mathcal{T}_{t-1,x,a},$ (2)
@@
$\Gamma_{t,i}:=\left[\sum_{\tau\in[t-1]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}-\mathcal{C}\right]_{0}^{\mathcal{C}},$
$\mathcal{C}\coloneqq 21L|X|\sqrt{2t|A|\ln\frac{2mT^{2}|X||A|}{\delta}}$ and $\cdot]_{a}^{b}\coloneqq\min(\max(\cdot,a),b)$. Finally, the weights are employed to build the estimates $\widehat{g}_{t,i}$ for each constraint $i\in[m]$ and each $(x,a)$ as the weighted mean of the values observed during the rounds in $\mathcal{T}_{t,x,a}$ (Line 9). Intuitively, the $\Gamma_{t}$ parameter allows the learning rates to meet the requirements of both the stochastic and the adversarial setting, as we point out in the following. In order to better understand it, we first introduce the following result.
@@
If $\beta_{t,i}(x,a)=\frac{1}{N_{t}(x,a)}$ for every $\tau\in\mathcal{T}_{t-1,x,a}$ such that $x_{\tau}=x,a_{\tau}=a$, then the following holds:
@@
$\widehat{g}_{t,i}(x,a)=\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t-1,x,a}}g_{\tau,i}(x,a).$
@@ 3.3 Decision Space Definition and Optimization Update
$\epsilon_{t}(x,a)=\sqrt{\frac{2|X_{k(x)+1}|\ln\frac{T|X||A|}{\delta}}{\max{1,N_{t}(x,a)}}},\ \forall(x,a)\in X\times A,$
@@
At each episode, the algorithm estimates the per-episode decision space $\Delta_{t}(\mathcal{P}_{t})$ taking the intersection between $\Delta(\mathcal{P}_{t})$ and the space of optimistically safe occupancy measures such that $(\widehat{g}_{t,i}-b_{t})^{\top}q\leq 0$ for all $i\in[m]$ (Line 11). We underline that the bonus quantity $b_{t}$ is necessary for the stochastic setting only, that is, when the empirical mean estimation is employed for the constraints. In the adversarial setting, the constraints estimator $\widehat{g}_{t,i}$ is sufficient to attain the desired theoretical guarantees.
Finally, the algorithm employs an online mirror descent (OMD) (Orabona, 2019) update step on the estimated feasible set $\Delta_{t}(\mathcal{P}_{t})$ (Line 12) employing the unnormalized Kullback-Leibler divergence as the Bregman divergence. Formally:
@@ 4.1 Stochastic Setting
In this section, we focus on the stochastic setting, that is, the constraints are sampled from fixed distributions. The first fundamental result is to show that the bonus terms $b_{t}(x,a)$ encompass the distance between the constraints estimator and the true constraint function. This is done by means of the following lemma.
@@
Intuitively, the result is proved as follows. We proceed by induction. In the first episodes, $\Gamma_{t,i}=0$ for all $i\in[m]$. Thus, by Proposition 3.1, the constraint estimator is computed as $\widehat{g}_{t,i}=\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t-1,x,a}}g_{\tau,i}(x,a),$ that is, the sample mean of the observed constraints values. Employing an Hoeffding bound, it is easy to see that Lemma 4.1 holds for those specific episodes. The induction step consists in showing that, assuming $\sum_{\tau\in[t-1]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}\leq\mathcal{C}$ at episode t1t-1, the same holds for the violation observed at $t$, too. This is done by showing that the empirical mean estimator and the bonus term are sufficient to keep the violation small when the constraints are stochastic. Again, since we proved that $\sum_{\tau\in[t]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}\leq\mathcal{C}$, we have $\Gamma_{t,i}=0$, which concludes the proof after a simple application of the Hoeffding inequality.
@@
Corollary 4.2 simply states that the true safe decision space is included in the per-episode decision space. This is intuitive, since, by Lemma 4.1, subtracting the bonus term to the constraints estimator allows, with high probability, to be optimistic in the constraints definition. A similar reasoning holds for the transitions. We are now ready to show the main result of the section, that is, the final regret and violation bound. This is done in the following theorem.
@@
$V_{t}\leq 61L|X|\sqrt{2t|A|\ln\left(\frac{2mT^{2}|X||A|}{\delta}\right)},\quad\forall t\in[T].$
Theorem 4.3 follows from the following reasoning. As concerns the regret bound, by Corollary 4.2, it holds that the safe optimum is included in the per-episode decision space, with high probability. Thus, following a standard no-regret argument of OMD with implicit exploration shows that Algorithm 1 attains sublinear regret with respect to any occupancy which is included in the algorithm decision space at each episode. Differently, to prove the violation, we proceed by contradiction, that is, we show that the weights definition does not allow the violation to exceed the threshold defined by the bound of Theorem 4.3. We remark that the proof for the violation is equivalent to the one for the adversarial setting, since the definition of $V_{t}$ is equivalent between the two settings. Indeed, in this case, we do not have to exploit Corollary 4.2, since even when $\Gamma_{t,i}=\mathcal{C}$, the violations are not allowed to exceed the aforementioned value.
@@
$\mathcal{V}_{t}\leq 18L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}},\quad\forall t\in[T].$
Intuitively, Theorem 4.4 is proved by showing that the positive violation attained by Algorithm 1 is proportional to the bonus $b_{t}$ term employed in the decision space definition. Showing that the term concentrate at a $1/\sqrt{T}$ rate concludes the proof. We finally remark that the results provided in this section strongly improve the ones provided in (Stradi et al., 2025d) for the stochastic setting, as we highlight in the following. First, Algorithm 1 does not rely on any Slater’s like condition to attain the optimal $\widetilde{\mathcal{O}}(\sqrt{T})$ regret and violation bounds. Second, Algorithm 1 attains the optimal rate for the positive constraints violation metric.
@@ 4.2 Adversarial Setting
$\alpha\text{-}R_{T}\leq 14L|X|^{2}\sqrt{2T|A|\ln\left(\frac{T|X|^{2}|A|}{\delta}\right)}$
@@ B.1 Results on the Optimization Update
Define $\tilde{q}_{t+1}$ such that:
$\tilde{q}_{t+1}(x,a,x^{\prime})=\widehat{q}_{t}(x,a,x^{\prime})e^{-\eta\widehat{\ell_{t}}(x,a)}.$
Since it holds:
$\widehat{q}_{t+1}=\operatorname*{argmin}_{q\in\Delta(\mathcal{P}_{t})}B(q||\tilde{q}_{t+1}),$
and
$\eta\widehat{\ell}_{t}^{\top}(\widehat{q}_{t}-q)=B(q\|\widehat{q}_{t})-B(q\|\tilde{q}_{t+1})+B(\widehat{q}_{t}\|\tilde{q}_{t+1}),$
by the condition $q\in\bigcap_{t\in[T]}\widehat{\Delta}_{t}(\mathcal{P}_{t})$ and the generalized Pythagorean theorem, it holds:
$B(q\|\widehat{q}_{t})\leq B(q\|\tilde{q}_{t+1}).$
Therefore, we have:
$\eta\sum_{t=1}^{T}\widehat{\ell}_{t}^{\top}(\widehat{q}_{t}-q)$ $=\sum_{t=1}^{T}B(q\|\widehat{q}_{t})-B(q\|\widehat{q}_{t+1})+B(\widehat{q}_{t}\|\tilde{q}_{t+1})$
$=B(q\|\widehat{q}_{1})-B(q\|\widehat{q}_{T+1})+\sum_{t=1}^{T}B(\widehat{q}_{t}\|\tilde{q}_{t+1})$
$\leq B(q\|\widehat{q}_{1})+\sum_{t=1}^{T}B(\widehat{q}_{t}\|\tilde{q}_{t+1}).$
To bound the term $B(\widehat{q}_{t}\|\tilde{q}_{t+1})$, we proceed as follows:
$B(\widehat{q}_{t}\|\tilde{q}_{t+1})$ $=\sum_{k=0}^{L-1}\sum_{x\in X_{k},a\in A,x^{\prime}\in X_{k+1}}\left(\eta\widehat{q}_{t}(x,a,x^{\prime})\widehat{\ell}_{t}(x,a)+\widehat{q}_{t}(x,a,x^{\prime})e^{-\eta\widehat{\ell}_{t}(x,a)}\right)$
$=\eta^{2}\sum_{k=0}^{L-1}\sum_{x\in X_{k},a\in A,x^{\prime}\in X_{k+1}}\widehat{q}_{t}(x,a,x^{\prime})\widehat{\ell}_{t}(x,a)^{2}$
$=\eta^{2}\sum_{x\in X,a\in A}\widehat{q}_{t}(x,a)\widehat{\ell}_{t}(x,a)^{2}$ (5)
$=\eta^{2}\sum_{x\in X,a\in A}\frac{\ell_{t}(x,a)}{u_{t}(x,a)+\gamma}\widehat{\ell}_{t}(x,a)$
$\leq\eta^{2}\sum_{x\in X,a\in A}\widehat{\ell}_{t}(x,a)$
$\leq\eta\sum_{x\in X,a\in A}\frac{q_{t}(x,a)}{u_{t}(x,a)}{\ell}_{t}(x,a)+\eta\frac{L\ln\frac{L}{\delta}}{2\gamma}$ (6)
$\leq\eta\sum_{x\in X,a\in A}{\ell}_{t}(x,a)+\eta\frac{L\ln\frac{L}{\delta}}{2\gamma}$
$\leq\eta|X||A|T+\eta\frac{L\ln\frac{L}{\delta}}{2\gamma},$ (7)
where Equation (5) is due to the fact that $1-e^{-z}\leq z$ for all $z\geq 0$, Inequality (6) holds with probability at least $1-\delta$ by Lemma 11 from [Jin et al., 2020] with $\alpha_{t}(x,a)=2\gamma$.
Choosing $q_{1}(x,a,x^{\prime})=\frac{1}{|X_{k}||A||X_{k+1}|}$ we have $B(q||q_{1})\leq L\ln(|X|^{2}|A|)$ and therefore:
$\eta\sum_{t=1}^{T}\widehat{\ell}_{t}^{\top}(\widehat{q}_{t}-q)$ $\leq\ln(|X|^{2}|A|)+\eta^{2}\sum_{x\in X,a\in A}q(x,a)\widehat{\ell}_{t}(x,a)^{2}$
$=L\ln(|X|^{2}|A|)+\eta^{2}|X||A|T+\eta^{2}\frac{L\ln\frac{L}{\delta}}{2\gamma}.$
Rearranging, we obtain the following final result:
$\sum_{t=1}^{T}\widehat{\ell}_{t}^{\top}(\widehat{q}_{t}-q)\leq L\frac{\ln(|X|^{2}|A|)}{\eta}+\eta|X||A|T+\eta\frac{L\ln\frac{L}{\delta}}{2\gamma}.$
This concludes the proof. ∎
@@ B.3 Results on the Weights
Consider a pair $(x,a)\in X\times A$, an index $i\in[m]$, and $t\in[T]$. By applying Lemma 5.3 of [Bernasconi et al., 2024], we obtain:
$w_{t,x,a,i}(\tau)$ $=\beta_{\tau,i}(x,a)\prod_{h=\tau+1}^{t-1}\left(1-\beta_{h,i}(x,a)\right)$
$=\frac{1}{N_{\tau}(x,a)}\prod_{h\in\mathcal{T}_{t-1,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right),\quad\forall\tau\in\mathcal{T}_{t-1,x,a}.$
We now focus on the term $\prod_{h\in\mathcal{T}_{t-1,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right)$:
$\prod_{h\in\mathcal{T}_{t-1,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right)$ $=\prod_{h\in\mathcal{T}_{t-1,x,a}:h>\tau}\frac{N_{h}(x,a)-1}{N_{h}(x,a)}$
$=\prod_{j=N_{\tau}(x,a)+1}^{N_{t-1}(x,a)}\frac{j-1}{j}$
$=\frac{N_{\tau}(x,a)}{N_{t-1}(x,a)}.$
@@
$w_{t,x,a,i}(\tau)=\frac{1}{N_{\tau}(x,a)}\frac{N_{\tau}(x,a)}{N_{t-1}(x,a)}=\frac{1}{N_{t-1}(x,a)}.$
@@
$V_{[t_{1},t_{2}],i}\leq\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}$ $\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau+1,i}(x,a)-\widehat{g}_{\tau,i}(x,a)\right)+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}$
$+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
@@
$\leq\sum_{\tau=t_{1}}^{t_{2}}g_{\tau,i}^{\top}q_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau,i}^{\top}\widehat{q}_{\tau}$ (12)
$=\sum_{\tau=t_{1}}^{t_{2}}g_{\tau,i}^{\top}q_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau}^{\top}\widehat{q}_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau,i}^{\top}\widehat{q}_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau,i}^{\top}q_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau,i}^{\top}q_{\tau}$

### v2

@@ 2.3 Baseline for the Adversarial Setting
where the unconstrained optimal value is defined as $\textsc{OPT}\coloneqq\max_{q\in\Delta(M)}\frac{1}{T}\sum_{t=1}^{T}r_{t}^{\top}q$. In order to quantify $\alpha$, we introduce a problem-specific parameter $\rho\in\left[0,1\right]$, which is defined as:
@@ 3 Algorithm
2:Initialize occupancy $\widehat{q}_{1}\leftarrow\frac{1}{|X_{k}||A||X_{k+1}|}$, the estimated transitions space $\mathcal{P}_{0}$ as the set of all the possible transition functions, and counters $N_{0}(x,a)=M_{0}(x^{\prime}|x,a)=0$ for all $k\in[0,...,L-1]$ and $(x,a,x^{\prime})\in X_{k}\times A\times X_{k+1}$
@@ 3.1 Initialization and Loss Estimation
Algorithm 1 receives as input the time horizon $T$, the set of states $X$, the set of actions $A$, the learning rate $\eta$, the implicit exploration factor $\gamma$, and the confidence $\delta\in(0,1)$. The occupancy measure $\widehat{q}_{1}$ is initialized uniformly over all tuples $(x_{k},a,x_{k+1})\in X_{k}\times A\times X_{k+1}$ for each layer $k\in[0,\dots,L-1]$. The transition function confidence set $\mathcal{P}_{1}$ is initialized as the set of all the possible transition functions. The counters $N_{t}(x,a)$ and $M_{t}(x^{\prime}|x,a)$, which are respectively defined as $N_{t}(x,a)=\sum_{\tau=1}^{t}\mathbb{I}_{\tau}\{x,a\}$ for all $(x,a)\in X\times A$, $M_{t}(x^{\prime}|x,a)=\sum_{\tau=1}^{t}\mathbb{I}_{t}\{x,a,x^{\prime}\}$ for all $(x,a,x^{\prime})\in X_{k}\times A\times X_{k+1},k\in[0,...,L-1],$ are initialized to 0 (see Line 2). We denote by $\mathbb{I}_{t}\{x,a\}$ and $\mathbb{I}_{t}\{x,a,x^{\prime}\}$ the indicator functions for the state-action(-state) visit at episode $\tau$.
@@ 3.2 Weights Estimation
At each episode, the counters are updated given the trajectory observed as feedback, namely, $N_{t}(x,a)$ and $M_{t}(x^{\prime}|x,a)$ are updated by incrementing by $1$ the entries of the tuples visited during the current episode. Then, the algorithm sets the weights that will be used to build the constraint estimates (Line 8). Specifically, given a pair $(x,a)\in X\times A$, $i\in[m]$, and $t\in[T]$, the weights $w_{t,x,a,i}$ are defined as follows:
$w_{t,x,a,i}(\tau)\coloneqq\beta_{\tau,i}(x,a)\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}(1-\beta_{h,i}(x,a))\quad\forall\tau\in\mathcal{T}_{t,x,a},$ (2)
@@
$\Gamma_{t,i}:=\left[\sum_{\tau\in[t]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}-\mathcal{C}_{t}\right]_{0}^{\mathcal{C}},$
$\mathcal{C}_{t}\coloneqq 21L|X|\sqrt{2t|A|\ln\frac{2mT^{2}|X||A|}{\delta}}$ and $\cdot]_{a}^{b}\coloneqq\min(\max(\cdot,a),b)$. Finally, the weights are employed to build the estimates $\widehat{g}_{t,i}$ for each constraint $i\in[m]$ and each $(x,a)$ as the weighted mean of the values observed during the rounds in $\mathcal{T}_{t,x,a}$ (Line 9). Intuitively, the $\Gamma_{t}$ parameter allows the learning rates to meet the requirements of both the stochastic and the adversarial setting, as we point out in the following. In order to better understand it, we first introduce the following result.
@@
If $\beta_{t,i}(x,a)=\frac{1}{N_{t}(x,a)}$ for every $\tau\in\mathcal{T}_{t,x,a}$, then the following holds:
@@
$\widehat{g}_{t,i}(x,a)=\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t,x,a}}g_{\tau,i}(x,a).$
@@ 3.3 Decision Space Definition and Optimization Update
$\epsilon_{t}(x,a)=\sqrt{\frac{2|X_{k(x)+1}|\ln\frac{T|X||A|}{\delta}}{\max\{1,N_{t}(x,a)\}}},\ \forall(x,a)\in X\times A,$
@@
At each episode, the algorithm estimates the per-episode decision space $\widehat{\Delta}_{t}(\mathcal{P}_{t})$ taking the intersection between $\Delta(\mathcal{P}_{t})$ and the space of optimistically safe occupancy measures such that $(\widehat{g}_{t,i}-b_{t})^{\top}q\leq 0$ for all $i\in[m]$ (Line 11). We underline that the bonus quantity $b_{t}$ is necessary for the stochastic setting only, that is, when the empirical mean estimation is employed for the constraints. In the adversarial setting, the constraints estimator $\widehat{g}_{t,i}$ is sufficient to attain the desired theoretical guarantees.
Finally, the algorithm employs an online mirror descent (OMD) (Orabona, 2019) update step on the estimated feasible set $\widehat{\Delta}_{t}(\mathcal{P}_{t})$ (Line 12) employing the unnormalized Kullback-Leibler divergence as the Bregman divergence. Formally:
@@ 4.1 Stochastic Setting
In this section, we focus on the stochastic setting, that is, the constraints are sampled from fixed distributions. The first fundamental result is to show that the bonus terms $b_{t}(x,a)$ encompass the distance between the constraint estimator and the true constraint function. This is done by means of the following lemma.
@@
Intuitively, the result is proved as follows. We proceed by induction. In the first episodes, $\Gamma_{t,i}=0$ for all $i\in[m]$. Thus, by Proposition 3.1, the constraint estimator is computed as $\widehat{g}_{t,i}=\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t,x,a}}g_{\tau,i}(x,a),$ that is, the sample mean of the observed constraint values. By employing an Hoeffding bound, it is easy to see that Lemma 4.1 holds for those specific episodes. The induction step consists in showing that, assuming $\sum_{\tau\in[t-1]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}\leq\mathcal{C}_{t-1}$ at episode t1t-1, the same holds for the violation observed at $t$. This is done by showing that the empirical mean estimator and the bonus term are sufficient to keep the violation small when the constraints are stochastic. Again, since we proved that $\sum_{\tau\in[t]}\sum_{x,a}g_{\tau,i}(x,a)\mathbb{I}_{\tau}\{x,a\}\leq\mathcal{C}_{t}$, we have $\Gamma_{t,i}=0$, which concludes the proof after a simple application of the Hoeffding inequality.
@@
Corollary 4.2 simply states that the true safe decision space is included in the per-episode decision space. This is intuitive, since, by Lemma 4.1, subtracting the bonus term to the constraint estimators allows, with high probability, to be optimistic in the constraints definition. A similar reasoning holds for the transitions. We are now ready to show the main result of the section, that is, the final regret and violation bound. This is done in the following theorem.
@@
$V_{t}\leq 61L|X|\sqrt{2t|A|\ln\left(\frac{2mT^{2}|X||A|}{\delta}\right)}\quad\forall t\in[T].$
Theorem 4.3 follows from the following reasoning. As concerns the regret bound, by Corollary 4.2, it holds that the safe optimum is included in the per-episode decision space, with high probability. Thus, a standard no-regret argument of OMD with implicit exploration shows that Algorithm 1 attains sublinear regret with respect to any occupancy that is included in the algorithm decision space at each episode. Differently, in order to prove the violation, we proceed by contradiction, that is, we show that the weights definition does not allow the violation to exceed the threshold defined by the bound of Theorem 4.3. We remark that the proof for the violation is equivalent to the one for the adversarial setting, since the definition of $V_{t}$ is equivalent between the two settings. Indeed, in this case, we do not have to exploit Corollary 4.2, since even when $\Gamma_{t,i}=\mathcal{C}_{t}$, the violations are not allowed to exceed the aforementioned value.
@@
$\mathcal{V}_{t}\leq 18L|X|\sqrt{2t|A|\ln\frac{2mT|X||A|}{\delta}}\quad\forall t\in[T].$
Intuitively, Theorem 4.4 is proved by showing that the positive violation attained by Algorithm 1 is proportional to the bonus $b_{t}$ term employed in the decision space definition. Showing that the term concentrate at a $1/\sqrt{T}$ rate concludes the proof. We finally remark that the results provided in this section strongly improve the ones provided in (Stradi et al., 2025d) for the stochastic setting, as we highlight in the following. First, Algorithm 1 does not rely on any Slater-like condition to attain the optimal $\widetilde{\mathcal{O}}(\sqrt{T})$ regret and violation bounds. Second, Algorithm 1 attains the optimal rate for the positive constraints violation metric.
@@ 4.2 Adversarial Setting
$\alpha\text{-}R_{T}\leq 14L|X|^{2}\sqrt{2T|A|\ln\left(\frac{T|X|^{2}|A|}{\delta}\right)},$
@@ B.1 Results on the Optimization Update
The result follows from Lemma 12 of [Jin et al., 2020], considering a general $q\in\bigcap_{t\in[T]}\widehat{\Delta}_{t}(\mathcal{P}_{t})$. ∎
@@ B.3 Results on the Weights
Consider a pair $(x,a)\in X\times A$, an index $i\in[m]$, and $t\in[T]$. By definition of the weights, it holds:
$w_{t,x,a,i}(\tau)$ $=\beta_{\tau,i}(x,a)\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}\left(1-\beta_{h,i}(x,a)\right)$
$=\frac{1}{N_{\tau}(x,a)}\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right),\quad\forall\tau\in\mathcal{T}_{t,x,a}.$
We now focus on the term $\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right)$:
$\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}\left(1-\frac{1}{N_{h}(x,a)}\right)$ $=\prod_{h\in\mathcal{T}_{t,x,a}:h>\tau}\frac{N_{h}(x,a)-1}{N_{h}(x,a)}$
$=\prod_{j=N_{\tau}(x,a)+1}^{N_{t}(x,a)}\frac{j-1}{j}$
$=\frac{N_{\tau}(x,a)}{N_{t}(x,a)}.$
@@
$w_{t,x,a,i}(\tau)=\frac{1}{N_{\tau}(x,a)}\frac{N_{\tau}(x,a)}{N_{t}(x,a)}=\frac{1}{N_{t}(x,a)}.$
@@
$V_{[t_{1},t_{2}],i}$ $\leq\sum_{x\in X,a\in A}\sum_{\tau\in\mathcal{T}_{t_{2},x,a}\cap[t_{1},t_{2}]}\frac{1}{\beta_{\tau,i}(x,a)}\left(\widehat{g}_{\tau,i}(x,a)-\widehat{g}_{\tau-1,i}(x,a)\right)+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}$
$\mkern 100.0mu+7L|X|\sqrt{2(t_{2}-t_{1})|A|\ln\frac{2T|X||A|}{\delta}},$
@@
$\leq\sum_{\tau=t_{1}}^{t_{2}}g_{\tau,i}^{\top}q_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau-1,i}^{\top}\widehat{q}_{\tau}$ (9)
$=\sum_{\tau=t_{1}}^{t_{2}}g_{\tau,i}^{\top}q_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}b_{\tau-1}^{\top}\widehat{q}_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau-1,i}^{\top}\widehat{q}_{\tau}+\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau-1,i}^{\top}q_{\tau}-\sum_{\tau=t_{1}}^{t_{2}}\widehat{g}_{\tau-1,i}^{\top}q_{\tau}$